### 22 Apr 2026 v3

# On Holomorphicity of Fiberwise Analytic Mappings

Abstract
We establish two rigidity results for mappings in complex geometry. First, we investigate the elevation of continuous mappings to holomorphic ones in the setting of proper fibrations. By leveraging relative morphism spaces and the Noetherian property of their local rings, we prove that a continuous, fiber-wise holomorphic map uniquely extends to a global holomorphic map, provided it is anchored by a sequence of holomorphic sections with mutually disjoint images. This result is then applied to establish the rigidity of homomorphisms between elliptic fibrations and Abelian schemes. Second, for a compact Kobayashi hyperbolic manifold fibered over $\mathbb{P}^{1}$, we prove that a continuous, fiber-wise holomorphic map of degree $1$ to a projective variety is a bi-holomorphic isomorphism, provided it is injective on a very ample hypersurface.
Keywords: Complex manifold, Holomorphic mapping, Relative morphism space, Elliptic fibration, Kobayashi hyperbolicity, Birational geometry
1 Introduction and Background
The rigid behavior of holomorphic functions of several complex variables is one of the foundational pillars of complex analysis. A striking manifestation of this rigidity is the ability to deduce global holomorphicity from merely partial and lower-dimensional analytic data. In 1906, F. Hartogs proved the following fundamental theorem on separate holomorphicity:
Theorem 1.1 (F. Hartogs, [5]).
Let $\Delta:=\{z\in\mathbb{C}:|z|<1\}$ be the open unit disc in the complex plane $\mathbb{C}$, and let $f\colon\Delta^{n}\rightarrow\mathbb{C}$ be an arbitrary function. Suppose that for every $(a_{1},\dots,a_{n})\in\Delta^{n}$ and every $i\in\{1,\dots,n\}$, the function
$g\colon\Delta\rightarrow\mathbb{C},\quad z\mapsto f(a_{1},\dots,a_{i-1},z,a_{i+1},\dots,a_{n})$
is holomorphic. Then $f$ itself is holomorphic.
In accordance with the spirit of rigidity exemplified by Hartogs’ theorem, this paper investigates how global structures in complex geometry can be enforced by partial algebraic constraints. We divide our exploration into two distinct but related results concerning the rigidity of mappings. In both settings, we demonstrate that the presence of a positive algebraic object, specifically an infinite sequence of disjoint sections or a very ample divisor, acts as a geometric anchor. When combined with appropriate local analytic conditions, this anchor elevates partial data to global holomorphic maps or global bi-holomorphic isomorphisms.
In Section 2, we turn our attention to the global extension of fiber-wise holomorphic mappings. While Hartogs’ theorem guarantees the extension of holomorphic functions across boundaries, elevating a continuous, fiber-wise holomorphic mapping to a globally holomorphic one on a proper fibration requires navigating the global geometry of the ambient space. We demonstrate that by anchoring the mapping to a sequence of holomorphic sections with mutually disjoint images, one can exploit the Noetherian nature of the local rings of the relative morphism space. This algebraic stabilization forces the mapping to be globally holomorphic, a rigidity result we apply to homomorphisms between elliptic fibrations and Abelian schemes.
In Section 3, we transition to the rigidity of proper holomorphic mappings. We present a rigidity theorem demonstrating how continuous, fiber-wise analytic mappings can be automatically elevated to global bi-holomorphic isomorphisms when anchored by a transverse algebraic intersection, specifically a very ample hypersurface, on a Kobayashi hyperbolic manifold.
2 Propagation of Holomorphicity on Proper Fibrations
We begin by investigating the conditions under which a continuous mapping, known to be holomorphic on the individual fibers of a proper fibration, can be elevated to a global holomorphic mapping. The strategy relies on translating the pointwise agreement of sections into an infinite intersection of complex analytic subsets within the relative morphism space. By invoking the Noetherian property of analytic local rings, this descending chain must stabilize, enforcing global holomorphicity.
Lemma 2.1.
Let $X,Y,Z$ be complex manifolds with $\dim(Y)=\dim(X)+1$ and $\phi\colon Y\rightarrow X$ a surjective proper holomorphic map with connected fibers. Let $\{s_{i}\}_{i\geq 1}$ be a sequence of holomorphic sections of $\phi$ with mutually disjoint images, and let $f\colon Y\rightarrow Z$ be a continuous mapping. Assume that $f|_{\phi^{-1}(x)}$ and $f\circ s_{i}$ are holomorphic for every $x\in X$ and all $i\geq 1$. Then $f$ is holomorphic.
Proof.
Let $U$ be a Zariski dense open subset of $X$ consisting of regular values of $\phi$. For any $x\in U$, the fiber $Y_{x}:=\phi^{-1}(x)$ is a compact Riemann surface. Since $Y_{x}$ is compact, the infinite set of points $\{s_{i}(x)\}_{i\geq 1}$ possesses an accumulation point in $Y_{x}$. By the analytic continuation principle for complex manifolds, any two holomorphic mappings from $Y_{x}$ to $Z$ that coincide on $\{s_{i}(x)\}_{i\geq 1}$ are identical.
Consider the relative morphism space
$\pi\colon\operatorname{Hom}_{U}(\phi^{-1}(U),Z\times U):=\{(x,g):g\in\operatorname{Hom}(Y_{x},Z),x\in U\}\rightarrow U,\quad(x,g)\mapsto x$
over $U$, of which connected components are complex analytic spaces by Douady’s theorem [2]. Let $\sigma\colon U\rightarrow\operatorname{Hom}_{U}(\phi^{-1}(U),Z\times U)$ be the continuous section defined by $\sigma(x):=(x,f_{x})$, and denote by $\mathcal{H}$ the connected component of $\operatorname{Hom}_{U}(\phi^{-1}(U),Z\times U)$ containing $\sigma(U)$.
For each integer $k\geq 1$, we define a holomorphic mapping $\Phi_{k}\colon\mathcal{H}\rightarrow(Z\times\cdots\times Z)\times U$ by
$\Phi_{k}(x,g):=(g(s_{1}(x)),\dots,g(s_{k}(x)),x),$
and define
$A_{k}:=\Phi_{k}^{-1}(\{(f(s_{1}(x)),\dots,f(s_{k}(x)),x):x\in U\}).$
Then $\{A_{k}\}_{k\geq 1}$ is a descending chain of complex analytic subsets in $\mathcal{H}$. By the analytic continuation principle, it holds that
$\bigcap_{k=1}^{\infty}A_{k}=\sigma(U).$
Because the local rings of $\mathcal{H}$ are Noetherian, the descending chain $\{A_{k}\}_{k\geq 1}$ must stabilize in a neighborhood of any point in $\sigma(U)$. Consequently, the restriction $\pi|_{\sigma(U)}$ is a bijective holomorphic mapping onto its image. Since $U$ is smooth, Zariski’s main theorem [4] immediately yields that $\sigma$ is holomorphic. By the universal property of the relative morphism space, the mapping $f|_{\phi^{-1}(U)}$ itself is holomorphic.
Since $f$ is continuous and the singular locus $Y-\phi^{-1}(U)$ is a proper analytic subset of $Y$, Riemann’s extension theorem applied locally in the coordinate charts of $Z$ guarantees that $f$ is holomorphic on the entirety of $Y$. ∎
The application of Lemma 2.1 converts algebraic data, namely the stabilization of descending analytic chains, into global analytic rigidity. We now apply this machinery to elevate continuous homomorphisms between elliptic fibrations and Abelian schemes into holomorphic morphisms. By utilizing the group law of the elliptic fibration, a single torsion-free holomorphic section generates the required sequence of disjoint sections.
Corollary 2.2.
Let $\pi\colon E\rightarrow B$ be a smooth elliptic fibration and $\varphi\colon A\rightarrow B$ an Abelian scheme over a complex manifold $B$. Let $f\colon E\rightarrow A$ be a continuous mapping such that $\pi=\varphi\circ f$ and
$f_{b}\colon\pi^{-1}(b)\rightarrow\varphi^{-1}(b),\quad P\mapsto f(P)$
is a homomorphism for every $b\in B$. Assume there exists a holomorphic section $\sigma$ of $\pi$ such that $f\circ\sigma$ is holomorphic and $\sigma(b)$ is torsion-free for every $b\in B$. Then $f$ is holomorphic.
Proof.
For each integer $k\geq 1$, we define a section $s_{k}\colon B\rightarrow E$ by $s_{k}(b):=k\sigma(b)$, where the multiplication is taken with respect to the group structure of the elliptic fibration $\pi$. Since $\sigma(b)$ is torsion-free for every $b\in B$, the sequence of sections $\{s_{k}\}_{k\geq 1}$ has mutually disjoint images.
Since $f$ is continuous, we have that $f_{b}$ is also continuous for all $b\in B$. But any continuous homomorphism between complex Lie groups is holomorphic, ensuring that $f|_{\pi^{-1}(b)}$ is holomorphic for all $b\in B$.
Because the restriction $f_{b}\colon\pi^{-1}(b)\rightarrow\varphi^{-1}(b)$ is a group homomorphism for every $b\in B$, we obtain
$(f\circ s_{k})(b)=f_{b}(k\sigma(b))=kf_{b}(\sigma(b))=k(f\circ\sigma)(b)$
for all $b\in B$. Therefore we have the identity $f\circ s_{k}=\mu_{k}\circ(f\circ\sigma)$, where $\mu_{k}\colon A\rightarrow A$ denotes the integer multiplication morphism on the smooth Abelian scheme $A$. Since $A$ is Abelian, the mapping $\mu_{k}$ is holomorphic. By the assumption that $f\circ\sigma$ is holomorphic, the composition $f\circ s_{k}$ is holomorphic for every integer $k\geq 1$.
The desired result then follows immediately from Lemma 2.1. ∎
Remark 2.3.
The proof of Corollary 2.2 encapsulates the overarching theme of this section: rigid global structures in complex geometry can be rigorously enforced by combining proper fibrations with the positive constraint of a torsion-free section. The algebraic stabilization within the relative morphism space serves as the essential bridge between continuous topological data and complex analytic rigidity.
3 Rigidity of Mappings from Kobayashi Hyperbolic Manifolds
We now transition from functions to mappings. In this section, we demonstrate that continuous maps from a Kobayashi hyperbolic manifold can be elevated to global isomorphisms, using a very ample hypersurface to lock the complex moduli.
Theorem 3.1.
Let $\phi\colon X\rightarrow Y$ be a continuous map of mapping degree $1$ between compact complex $n$-manifolds of equal dimension $n\geq 4$, where $Y$ is projective and $X$ is Kobayashi hyperbolic. Suppose that there exist a holomorphic fibration $f\colon X\rightarrow\mathbb{P}^{1}$ and a very ample hypersurface $H$ in $X$ such that $\phi|_{f^{-1}(t)}$ is holomorphic for every $t\in\mathbb{P}^{1}$ and $\phi|_{H}$ is injective. Then $\phi$ is a bi-holomorphic isomorphism.
Proof.
We divide the proof into three main steps: Establishing that the restrictions of $\phi$ on the fibers are finite, then proving their images are mutually disjoint, and finally upgrading the continuous map $\phi$ to a biholomorphism.
Step 1: Since $H\in\operatorname{div}(X)$ is ample, the complex manifold $X$ is projective by Kodaira’s embedding theorem [3]. Let
$\Sigma:=\mathbb{P}^{1}-\{b_{1},\dots,b_{m}\}$
be the punctured Riemann sphere consisting of the regular values of $f$. For any $t\in\Sigma$, the fiber $X^{t}:=f^{-1}(t)$ is a smooth projective subvariety of $X$ of pure dimension n1n-1. Since $\phi^{t}:=\phi|_{X^{t}}$ is holomorphic and $Y$ is projective, Remmert’s proper mapping [4] theorem and Chow’s theorem [3] ensure that the image $Y^{t}:=\phi(X^{t})$ is a projective subvariety of $Y$. We first establish the injectivity of the restriction $\phi^{t}$ for any fixed $t\in\Sigma$. For the sake of contradiction, we now assume that $\phi^{t}$ is not injective.
Suppose first that $\dim(Y^{t})\leq n-3$. By Chevalley’s fiber dimension theorem, the generic fiber of $\phi^{t}$ would have dimension at least 2, and hence would contain a projective surface. Since the intersection of the ample divisor $H$ with a projective surface contains a curve in $H$, the morphism $\phi^{t}$ would contract a curve in $H$ to a point, violating the injectivity of $\phi|_{H}$. Therefore, we must have at least $\dim(Y^{t})\geq n-2$.
Suppose now that $\dim(Y^{t})=n-2$. Take a generic point $y\in Y^{t}$. Then the fiber $C$ of $\phi^{t}$ over $y$ is a projective curve. Since $H\in\operatorname{div}(X)$ is a very ample divisor, $H^{t}:=X^{t}\cap H$ is very ample in $X^{t}$, making $H^{t}\cap C$ non-empty and $\phi(H^{t})=Y^{t}$. Because $\phi|_{H}$ is injective by assumption, the mapping degree of $\phi^{t}|_{H^{t}}$ onto its image is exactly 1. By choosing $y$ to be generic, we ensure that the restriction $\phi^{t}|_{H^{t}}$ is unramified at its unique preimage $x\in H^{t}$, rendering its differential an isomorphism at $x$. As a result, the tangent space $T_{x}H^{t}$ intersects trivially with the kernel of the differential of $\phi^{t}$ at $x$, which is precisely $T_{x}C$. This proves that $C$ intersects $H^{t}$ transversally, and hence their intersection product is $H^{t}\cdot C=1$. Because $H$ is a very ample divisor, intersecting a projective curve $C$ at a single point with a multiplicity of exactly 1 forces $C$ to be rational, that is, bi-holomorphic to $\mathbb{P}^{1}$. However, as $X$ is Kobayashi hyperbolic, there exists no non-constant holomorphic mapping from $\mathbb{P}^{1}$ into $X$ by Brody’s theorem [1] on hyperbolicity. This is a contradiction. Therefore, we have in fact $\dim(Y^{t})=n-1$, so the regular mapping $\phi^{t}\colon X^{t}\rightarrow Y$ sends its domain of definition onto a projective variety $Y^{t}$ of equal dimension n1n-1.
We claim that $\phi^{t}$ is actually a finite morphism. Consider the Stein factorization $\phi^{t}=\phi^{t}_{1}\circ\phi^{t}_{2}$ of $\phi^{t}$, where $\phi^{t}_{2}\colon X^{t}\rightarrow N$ is a surjective projective morphism with connected fibers onto a normal projective variety $N$, and $\phi^{t}_{1}\colon N\rightarrow Y^{t}$ is a finite morphism. Since $\phi^{t}_{1}$ is finite and surjective, we obtain
$\dim(N)=\dim(Y^{t})=n-1.$
Because $\phi^{t}_{2}$ is a surjective morphism with connected fibers between projective varieties of the same dimension, its generic fiber is a single point, implying that $\phi^{t}_{2}$ is a birational morphism. Assume to the contrary that $\phi^{t}$ admits a fiber $F$ of positive dimension. Since $\phi^{t}_{1}$ is finite, this fiber $F$ must be contracted to a point entirely by $\phi^{t}_{2}$. Therefore, the exceptional locus of the birational morphism $\phi^{t}_{2}$ is non-empty. Now, recall from birational geometry that the exceptional locus of a proper birational morphism from a smooth projective variety is uniruled, meaning it is covered by rational curves. However, as $X$ is Kobayashi hyperbolic, by Brody’s theorem [1], it contains no rational curves, which leads to a contradiction. This proves that $\phi^{t}$ is indeed a finite morphism.
Step 2: Now, we prove that $Y^{t_{1}}$ and $Y^{t_{2}}$ cannot intersect for distinct $t_{1},t_{2}\in\Sigma$. Assume to the contrary that there exist points $x_{1}\in X^{t_{1}}$ and $x_{2}\in X^{t_{2}}$ for distinct $t_{1},t_{2}\in\Sigma$ such that $\phi(x_{1})=\phi(x_{2})$. Since $\phi^{t}$ is a finite surjective morphism onto its image for every $t\in\Sigma$, the images $Y^{t_{1}}$ and $Y^{t_{2}}$ are projective subvarieties of $Y$ of dimension n1n-1. By Krull’s dimension theorem, the intersection $Y^{t_{1}}\cap Y^{t_{2}}$ is a projective subvariety of $Y$ with dimension bounded below by
$\dim(Y^{t_{1}}\cap Y^{t_{2}})\geq(n-1)+(n-1)-n=n-2\geq 2.$
Consider the incidence variety
$I:=\{(a_{1},a_{2})\in X^{t_{1}}\times X^{t_{2}}\mid\phi(a_{1})=\phi(a_{2})\}.$
Because $\phi^{t_{1}}$ and $\phi^{t_{2}}$ are finite morphisms, the natural projection from $I$ onto $Y^{t_{1}}\cap Y^{t_{2}}$ has finite fibers. Therefore, $I$ is a projective subvariety of the fiber product $X^{t_{1}}\times X^{t_{2}}$ with
$\dim(I)=\dim(Y^{t_{1}}\cap Y^{t_{2}})\geq 2.$
Let $p_{1}\colon I\rightarrow X^{t_{1}}$ and $p_{2}\colon I\rightarrow X^{t_{2}}$ be the restrictions of the canonical projections onto each factor. Because $\phi^{t_{1}}$ and $\phi^{t_{2}}$ are finite morphisms, the projections $p_{1}$ and $p_{2}$ are also finite. Therefore, the pullbacks $p_{1}^{*}(H)$ and $p_{2}^{*}(H)$ are both ample divisors on the projective variety $I$. Since $\dim(I)\geq 2$, their intersection product $p_{1}^{*}(H)\cdot p_{2}^{*}(H)$ is strictly positive, ensuring the intersection of their support is non-empty. Take any $(h_{1},h_{2})\in p_{1}^{*}(H)\cap p_{2}^{*}(H)$. Then, by the very construction, we have $h_{1},h_{2}\in H$ such that $\phi(h_{1})=\phi(h_{2})$, which again contradicts the injectivity of $\phi|_{H}$. We conclude that the images $Y^{t}$ are mutually disjoint for all $t\in\Sigma$.
Step 3: Because $Y$ is a smooth projective variety, and the collection $\{Y^{t}:t\in\Sigma\}$ consists of homologous effective divisors that are pairwise disjoint, the intersection product $Y^{t_{1}}\cdot Y^{t_{2}}$ vanishes for all distinct $t_{1},t_{2}\in\Sigma$. Therefore, a Zariski dense open subset of the compact complex manifold $Y$ is continuously foliated by the hypersurfaces $Y^{t}$ over $\Sigma$, yielding a continuous mapping
$\upsilon\colon\Sigma\rightarrow\operatorname{Chow}(Y),\quad t\mapsto Y^{t}$
into the Chow variety of $Y$. Because the connected components of $\operatorname{Chow}(Y)$ are all projective, we may consider the Zariski closure $V$ of the image $\upsilon(\Sigma)$ within $\operatorname{Chow}(Y)$. Since $\Sigma$ is an uncountably infinite set and the members of $\upsilon(\Sigma)$ are mutually disjoint, the image $\upsilon(\Sigma)$ consists of uncountably many distinct points. As a 0-dimensional variety over $\mathbb{C}$ is necessarily a finite set of points, the Zariski closure $V$ must have dimension $\dim(V)\geq 1$. This implies that $V$ is a projective family of effective divisors on $Y$ of strictly positive dimension. Because the condition of having empty intersections is an open condition in the Zariski topology, and $\upsilon(\Sigma)$ is by construction Zariski dense in $V$, the generic members of the projective family $V$ are also mutually disjoint. Since the effective divisors in $V$ are algebraically equivalent and generically disjoint, their intersection product vanishes. On any smooth projective variety, a positive-dimensional projective family of effective divisors with vanishing mutual intersection numbers must possess an empty base locus. Consequently, the linear equivalence classes of divisors in $V$ span a free pencil, which guarantees the existence of a surjective regular morphism $g\colon Y\rightarrow\mathbb{P}^{1}$ such that $g^{-1}(t)=Y^{t}$ for every $t\in\mathbb{P}^{1}$. Because $\phi$ maps the fibers of $f$ into the fibers of $g$, we have $g\circ\phi=f$. Since the mapping degree is multiplicative, we have
$\deg(\phi^{t})=\deg(\operatorname{id}_{\mathbb{P}^{1}})\cdot\deg(\phi^{t})=\deg(\phi)=1$
for all $t\in\mathbb{P}^{1}$. Since $Y$ is a smooth projective variety and $g\colon Y\rightarrow\mathbb{P}^{1}$ is a holomorphic fibration, Bertini’s theorem [6] ensures that the general fiber of $g$ is smooth, and therefore normal. Thus, for a generic $t\in\Sigma$, the finite morphism $\phi^{t}\colon X^{t}\rightarrow Y^{t}$ of degree 1 onto a normal variety is an isomorphism by Zariski’s main theorem [6]. This in particular proves that $\phi$ is injective almost everywhere.
Take an arbitrary $t\in\Sigma$. Since $X$ is a compact Kobayashi hyperbolic manifold, we have that $X^{t}$ is also Kobayashi hyperbolic. Recall from birational geometry that any rational map into a projective variety containing no rational curves is automatically a regular morphism. Therefore, the group of birational mappings from $X^{t}$ to itself coincides with $\operatorname{Aut}(X^{t})$, which is finite as
$H^{0}(X^{t};\operatorname{Der}(\mathcal{O}_{X^{t}}))=0$
by the rigidity of Kobayashi hyperbolic manifolds. Consider the relative morphism scheme $\Phi\colon\operatorname{Hom}_{\mathbb{P}^{1}}(X,Y)\rightarrow\mathbb{P}^{1}$, where $X$ and $Y$ are viewed as fibrations over $\mathbb{P}^{1}$ via $f$ and $g$ respectively. For any two finite morphisms $\varphi_{1},\varphi_{2}\colon X^{t}\rightarrow Y^{t}$ of degree 1, the composition
$(\varphi_{1})^{-1}\circ\varphi_{2}\colon X^{t}\rightarrow X^{t}$
is birational. Consequently, the set of all finite morphisms from $X^{t}$ to $Y^{t}$ of degree 1 forms a principal homogeneous space under the $\operatorname{Aut}(X^{t})$-action. Since $\operatorname{Aut}(X^{t})$ is finite, the set $Z^{t}$ of all finite morphisms of degree 1 in the fiber of $\Phi\colon\operatorname{Hom}_{\mathbb{P}^{1}}(X,Y)\rightarrow\mathbb{P}^{1}$ over any $t\in\Sigma$ is finite. Moreover, since the Zariski closure of
$Z:=\{(t,\varphi):t\in\Sigma,\varphi\in Z^{t}\}$
in $\operatorname{Hom}_{\mathbb{P}^{1}}(X,Y)$ is precisely the connected component of $\operatorname{Hom}_{\mathbb{P}^{1}}(X,Y)$ containing the finite morphism $\phi^{t}$ of mapping degree 1, it is a projective curve such that $\Phi|_{Z}$ is a branched covering. Since $\phi$ is uniformly continuous, it induces a continuous section
$\phi^{*}\colon\Sigma\rightarrow Z,\quad t\mapsto\phi^{t}$
of the branched covering $\Phi|_{Z}$. Since $\Phi|_{Z}$ admits a continuous section over $\Sigma$, it is in fact not branched over $\Sigma$. Moreover, any continuous section of an unbranched analytic covering is necessarily holomorphic, so $\phi^{*}$ is holomorphic on $\Sigma$. Because $\phi$ is holomorphic along fibers and the section $\phi^{*}$ is also holomorphic, Osgood’s lemma [4] on separate holomorphicity guarantees that $\phi$ is holomorphic on $f^{-1}(\Sigma)$. Since $\phi$ is continuous and the singular locus $X-f^{-1}(\Sigma)$ forms a proper algebraic subset, Riemann’s extension theorem guarantees that $\phi$ is holomorphic on the entirety of $X$.
Having established that $\phi\colon X\rightarrow Y$ is a proper holomorphic mapping of degree 1 between smooth projective varieties, it remains to show it is an isomorphism. But, if $\phi$ were not an isomorphism, as it is injective almost everywhere, it would be a birational map contracting an exceptional divisor to a lower-dimensional subvariety. Recall from birational geometry that the exceptional locus of a proper birational morphism between smooth projective varieties is uniruled, which once again contradicts the hyperbolicity of $X$.
Quod erat demonstrandum. ∎
Remark 3.2.
Theorem 3.1 illustrates another facet of mapping regularity in complex geometry. While the results in Section 2 rely on an infinite sequence of disjoint sections to propagate holomorphicity, this theorem demonstrates that a single very ample hypersurface provides sufficient algebraic anchoring to elevate a continuous mapping to a global bi-holomorphic isomorphism. The intrinsic absence of rational curves in the Kobayashi hyperbolic manifold prevents the formation of exceptional loci, ensuring that the induced birational morphism is necessarily an isomorphism.
Acknowledgement
The author is deeply grateful to Weiyi Zhang for very useful and inspiring discussions. The author expresses gratitude to the reviewers for various suggestions.
This research was completed while the author was studying at the Mathematics Institute of the University of Warwick. The author therefore would like to thank the University of Warwick for its hospitality.
The author is in debt to the Russian mathematical society, especially, to all the professors teaching the math in Moscow program, and most importantly, to the author’s supervisor Alexander Petrovich Veselov at Loughborough University, as they cultivated the author’s mathematical literacy and maturity.
Funding Information
No funding was received to assist with the preparation of this manuscript, and the author did not receive support from any organization for the submitted work.
Statements and Declarations
The author certifies that the author has no affiliations with or involvement in any other organization or entity with any financial interest or non-financial interest in the subject matter or materials discussed in this manuscript. This article is licensed under a Creative Commons Attribution 4.0 International License, which permits use, sharing, adaptation, distribution and reproduction in any medium or format, as long as you give appropriate credit to the original author and the source, provide a link to the Creative Commons licence, and indicate if changes were made. The images or other third party material in this article are included in the article’s Creative Commons licence, unless indicated otherwise in a credit line to the material. If material is not included in the article’s Creative Commons licence and your intended use is not permitted by statutory regulation or exceeds the permitted use, you will need to obtain permission directly from the copyright holder.
Data sharing is not applicable to this article as no datasets were generated or analysed during the current study.
The author hereby provides consent for the publication of the manuscript detailed above.